Rizza's age is one year more than twice Norie's age.Eight years ago,the sum of their ages was equal to Rizza's age ten years from now.How old are they now?
Answer by gwendolyn(128)

Let Rizza's age=R
Let Norie's age=N
Let Rizza's age 8 years ago=R-8
Let Norie's age 8 years ago=N-8
Let Rizza's age 10 years from now=N+10
Two equations can be formed from the information in the question and the variables I have set.
R=2N+1
(R-8)+(N-8)=R+10
First, we cancel out the two 'R's in the second equation:
-8+N-8=10
Then we should add 16 to both sides.
N=26
Next, we should plug in our value of N.
R=2N+1
R=54+1
R=55
So, Rizza is 55 and Norie is 26.